# Attitudes Towards and Knowledge of Non-Consensual Synthetic Intimate Imagery in 10 Countries

### 26 Jan 2024 v1

@@ 3.1. Ethical Considerations
In mid-2023, we conducted a cross-country online survey on image-based sexual abuse, defined as the non-consensual taking, creating, or sharing of intimate images (photos or videos), including threats to share intimate images. The study procedure and survey were approved by the Human Research Ethics Committee at RMIT University. Respondents were informed of the topic at the beginning of the survey and were allowed to opt out at any time. All questions involving self-reporting of victimization and/or perpetration were voluntary and respondents could choose an answer of ”prefer not to say” while still completing the survey. Respondents were given support service information tailored to their geographic location at the end of the survey. Respondents were compensated based on pilot tests of time taken in each country and exchange rates.
@@ 3.2. Study Design
We conducted the online survey in ten countries, employing a research and survey firm (YouGov) to survey a minimum of 1,600 adults (18+) in each country, for a total respondent count of 16,693. YouGov was selected due to their comprehensive country coverage and large panels (Sandle, 2021), which should reduce the likelihood of needing extensive within-country weighting. YouGov retains demographic information on their survey panelists, and screener questions were then used to achieve representative sampling based on quotas. Upon qualifying for the survey, completion rates ranged from 71.0% (Belgium) to 89% (Australia). The final dataset consisted of respondents who passed the YouGov standard quality checks, including bot catchers, speeder detection, and manual checking of open-text questions. In conducting our survey, we intentionally chose a variety of countries which represent the spectrum of how legislation has dealt with NSII, image-based abuse, and pornography more broadly. The goal was to obtain representative samples by age, gender, and location (e.g., state or territory in Australia, province in Korea, etc) within each country, using the latest official population estimates reflecting each surveyed country. The survey built on an earlier survey conducted in Australia (Powell and Henry, 2019), which focused on non-consensual intimate imagery more broadly (but did not specifically ask about deepfakes). The revised survey (median completion time 18.2 minutes) added some additional scales, teased apart how content was obtained (e.g., filmed/photographed versus stolen from a device/cloud), and included questions on NSII specifically.
In this paper, we focus on the subset of 23 questions in our survey relating to NSII and digitally altered intimate images. To administer the survey in the official language(s) of each country, the survey vendor provided translations into French, Spanish, Danish, Dutch, Polish, and South Korean. These were then double-checked and edited as needed by native speaker colleagues of the research team for quality assurance. We aimed to have as similar language as possible across countries, while also accounting for cultural differences in comfort speaking about sexual topics. The survey was deployed from May to June 2023. In some countries, samples were not fully representative. Accordingly, when presenting dis-aggregated findings, the data are weighted by age, gender, and location. When we aggregate across countries, the data are additionally weighted by population, rounded to the nearest 1,000. The weighted statistics are presented throught the quantitative section, and all quantitative analyses were performed using R Statistical Software (v4.3.2; R Core Team 2023)(R Core Team, 2023). Statistics presented in the qualitative section are unweighted. We primarily used the survey (Lumley, 2023), srvyr (Freedman Ellis and Schneider, 2023), marginaleffects (arel2023marginal), and epitools (Aragon, 2020) packages for statistical analyses. When calculating confidence intervals for proportions, we use the Rao-Scott scaled chi-squared distribution for the loglikelihood from a binomial distribution (e.g., % of respondents who had experienced victimization). When calculating within-subject mean differences, we use two tailed one-sample weighted t-tests with bootstrapped standard errors, using 1,000 replicates (e.g., differences in criminalization attitudes across different types of behaviors). When calculating between-gender mean differences we use two tailed two-sample weighted Welch t-tests using bootstrapped standard errors, using 1,000 replicates (e.g., difference in how much women think sharing deepfake pornography of celebrities should be criminalized as compared to men). We used the svyglm function with a quasibinomial family to calculate risk ratios for binary categorical values (e.g. comparing how much more likely men were to report having watched deepfake pornography as compared to women). When evaluating differences amongst categorical variables in the unweighted qualitative data, we use Pearson’s Chi-squared test with Yates’ continuity correction.
@@
The mean respondent age was 46.0 years ($sd=16.7$). Women accounted for 50.9% of the respondents, men for 47.6%, and together ”other,” ”prefer not to say,” and ”non-binary” accounted for the remaining 1.4%. Detailed demographic breakdowns by country are available in Table A1 in the Appendix. Our gender results and statistical tests focus primarily on the users who identified as women or men, excluding the 1.0% of other users for two reasons: we are using census weights, which are often limited to binary genders, and our total sample (unweighted) of non-binary respondents was 93, which limits our ability to perform statistical tests for significance.
@@ 3.3. Measures and Analysis
In relation to NSII, all respondents were asked three closed-ended questions, as well as one optional open-ended question presented towards the end of the survey. They were first asked whether they were familiar with the concept of deepfake pornography using the following wording: ”Realistic-looking fake porn can now be created using artificial intelligence (AI) to swap the faces of pornographic actors with other people’s faces, so that it looks like they’re in a porn video. This is sometimes referred to as ’deepfake pornography’ as it looks very realistic and can be hard to recognize as fake or digitally created. How familiar are you with this concept?”. This question offered four possible categorical response options, as seen in Table 1. Second, respondents were asked how much they thought a series of behaviors were worthy of criminalization (behaviors can be seen in Figure 1). Response options ranged from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with the midpoint of 0 being ”Not sure.” And third, respondents were asked how much they agreed or disagreed with the following statement: ”People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ‘deepfake’) of them without their permission.” Respondents answered this question on a 7-point Likert scale, anchored at 1=Strongly disagree and 7=Strongly agree (we present the mean alongside the median in the results - see below) (Sullivan and Artino Jr, 2013).
@@
The remaining close-ended questions were asked of a subset of respondents, depending on their knowledge and experiences. A subset of respondents who indicated at least some familiarity (i.e., ”I know a little bit about it” or ”I am quite familiar with this”) were asked follow-up questions around NSII behaviors as seen in Figure 2. As part of a broader set of questions regarding the victimization and perpetration of image-based sexual abuse since the age of 18, we also asked the subset of respondents who indicated a previous experience with digitally altered images whether the relevant content was a video, and if so, whether the video was a deepfake (allowing respondents to select yes/no/don’t know, or prefer not to answer). For victimization questions, the wording was intentionally not blame-casting. Similarly, perpetration questions were phrased neutrally to focus on whether the respondent had ever done an action.
@@
All respondents could opt into answering the open-ended question with the following wording: ”Would you like to add any additional thoughts on the topic of deepfakes? Please feel free to skip this question if you don’t have anything to add.” We analyzed the qualitative data for the open-ended response questions using an iterative, inductive approach. Our approach was to translate the responses from origin language to English via the Google Translate API , familiarize ourselves with the data, develop a codebook, code the data, and then review the data to make sense of the themes from the first coding process. To do so, two members of the research team read through 25% of the sample responses across the countries and then co-developed a codebook that detailed descriptions and restrictions on what could be included within a code and provided concrete examples of each code. The same two members of the research team coded the first 150 responses based on four preliminary codes that emerged from the initial review of the data. Coder 1 and coder 2 independently moved through the subset of data labelling the open-ended responses with relevant codes. Following this, coder 1 and 2 compared coding allocations for quality assurance. The two coders then identified and removed responses that were deemed inappropriate or irrelevant to the research (e.g., ”no comment,” or ”no thanks, I do not know this topic”). Finally, the team revised and discussed the themes that emerged from the responses (n = 2,474). The final codebook contained four key themes: ”condemnation,” ”punitive and justice attitudes,” ”a fear of technology,” and ”victim-blaming.” Interrater reliability (IRR) for each code was calculated using Cohen’s Kappa. Across the four codes, Kappa values ranged from .135 to .225, with a median of .173.
@@ 4. Results
Results are organized into four sections. The first section addresses the first research question, focusing on awareness and attitude questions that were asked of all respondents. The second section answers the second question by focusing on behaviors involving NSII, including victimization, perpetration, and consumption. The third section investigates the role of gender in influencing answers to the first and second research questions. Finally, the last section is a qualitative analysis of the open-ended opinion question. We present our results primarily descriptively and holistically, particularly at the country level, as we have no reason to a priori assume differences between countries, given the lack of prior data. When discussing gender differences, statistical analyses (accounting for weights) are presented, alongside measures of uncertainty. We conclude with some qualitative trends from the single, open-end question to which respondents could choose to respond.
@@ 4.1. Awareness and Attitudes
Country Response % of Respondents 95% CI
Australia I have never heard of this until now 36.8 35.4-38.2
I have only heard the phrase 30.3 29-31.7
I know a little bit about it 24.7 23.4-26
I am quite familiar with this 8.2 7.3-9.1
Belgium I have never heard of this until now 40.1 39.1-41.1
I have only heard the phrase 28.0 27.1-28.9
I know a little bit about it 25.6 24.7-26.5
I am quite familiar with this 6.3 5.8-6.8
Denmark I have never heard of this until now 34.8 34.1-35.5
I have only heard the phrase 44.1 43.4-44.8
I know a little bit about it 17.0 16.5-17.6
I am quite familiar with this 4.10 3.8-4.4
France I have never heard of this until now 55.2 52.8-57.6
I have only heard the phrase 21.6 19.7-23.6
I know a little bit about it 18.4 16.6-20.3
I am quite familiar with this 4.8 3.8-5.9
Mexico I have never heard of this until now 51.7 48.8-54.7
I have only heard the phrase 22.3 19.9-24.8
I know a little bit about it 21.0 18.7-23.5
I am quite familiar with this 4.9 3.8-6.3
Netherlands I have never heard of this until now 29.0 27.9-30.1
I have only heard the phrase 31.7 30.6-32.8
I know a little bit about it 31.8 30.7-32.9
I am quite familiar with this 7.5 6.9-8.2
Poland I have never heard of this until now 50.9 49.1-52.8
I have only heard the phrase 30.1 28.4-31.8
I know a little bit about it 15.2 13.9-16.5
I am quite familiar with this 3.8 3.2-4.6
South Korea I have never heard of this until now 31.1 29.1-33.2
I have only heard the phrase 39.2 37.1-41.4
I know a little bit about it 26.6 24.7-28.6
I am quite familiar with this 3.0 2.3-3.8
Spain I have never heard of this until now 49.5 47.5-51.5
I have only heard the phrase 27.6 25.8-29.4
I know a little bit about it 19.5 17.9-21.1
I am quite familiar with this 3.5 2.8-4.3
USA I have never heard of this until now 42.5 37.7-47.4
I have only heard the phrase 26.9 22.7-31.4
I know a little bit about it 23.7 19.6-28.1
I am quite familiar with this 6.9 4.5-9.8
@@
Figure 1. Heatmap of Mean Criminalization Attitudes by Country. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with a midpoint of 0 (”Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parentheses
@@
[Heat map with country on the x axis and experience type on the y axis]Heat map with country on the x axis and experience type on the y axis. Mean weighted scores are presented alongside standard errors in parantheses. All scores are positive, indicating that on average, respondents fell on the side of ”should be a crime”.
@@
Figure 1. Heatmap of Mean Criminalization Attitudes by Country. Scale ranges from -2 (”Definitely should not be a crime”) to 2 (”Definitely should be a crime”), with a midpoint of 0 (”Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parentheses
[…]
@@
Finally, respondents were asked about how much they agreed or disagreed with the following statement: ”People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ’deepfake’) of them without their permission,” with anchors of 1=Strongly disagree and 7=Strongly agree. As seen in Table 4.1, across the different countries, the mean response fell between 1 and 2, with a median of 1 in all countries, indicating a general sentiment that people should get upset if they are the target of a deepfake pornography video.
@@
Table 2. Level of agreement with statement: ”People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ‘deepfake’) of them without their permission” (1 = strongly disagree, 4 = neither agree not disagree, 7=strongly agree). Weighted statistics presented.

### 13 Feb 2024 v2

@@ 3.1. Ethical Considerations
In mid-2023, we conducted a multi-country online survey on image-based sexual abuse, defined as the non-consensual taking, creating, or sharing of intimate images (photos or videos), including threats to share intimate images. The study procedure and survey were approved by the Human Research Ethics Committee at the Royal Melbourne Institute of Technology. Respondents were informed of the topic at the beginning of the survey and were allowed to opt out at any time. All questions involving self-reporting of victimization and/or perpetration were voluntary and respondents could choose an answer of “prefer not to say” while still completing the survey. Respondents were given support service information tailored to their geographic location at the end of the survey. Respondents were compensated based on pilot tests of time taken in each country and exchange rates.
@@ 3.2. Study Design
We conducted the online survey in 10 countries, employing a research and survey firm (YouGov) to survey a minimum of 1,600 adults (18+) in each country, for a total respondent count of 16,693. YouGov was selected due to their comprehensive country coverage and large panels (Sandle, 2021), which reduces the likelihood of needing extensive within-country weighting. YouGov retains demographic information on their survey panelists, and screener questions were used to achieve representative sampling based on quotas. Upon qualifying for the survey, completion rates ranged from 71.0% (Belgium) to 89% (Australia). The final dataset consisted of respondents who passed the YouGov standard quality checks, including bot catchers, speeder detection, and manual checking of open-text questions.
In conducting our survey, we intentionally chose a variety of countries that represent the spectrum of how legislation has dealt with NSII, image-based abuse, and pornography more broadly. The goal was to obtain representative samples by age, gender, and location (e.g., state or territory in Australia, province in South Korea, etc.) within each country, using the latest official population estimates reflecting each surveyed country. The survey built on an earlier survey conducted in Australia (Powell and Henry, 2019), which focused on non-consensual intimate imagery more broadly (but did not specifically ask about deepfakes). The revised survey (median completion time: 18.2 minutes) added some additional scales, teased apart how content was obtained (e.g., filmed/photographed versus stolen from a device/cloud), and included questions on NSII and “deepfake pornography” specifically.
In this paper, we focus on the subset of 23 questions in our survey relating to AI-generated image-based sexual abuse (AI-IBSA). To administer the survey in the official language(s) of each country, the survey vendor provided translations into Danish, Dutch, French, Polish, Spanish, and South Korean. These were then double-checked and edited as needed by native speaker colleagues of the research team for quality assurance. We aimed to have as similar language as possible across countries, while also accounting for cultural differences in comfort speaking about sexual topics.
The survey was deployed from May to June 2023. In some countries, samples were not fully representative. Accordingly, when presenting disaggregated findings in the quantitative section, the data are weighted by age, gender, and location. When we aggregate across countries, the data are additionally weighted by population, rounded to the nearest 1,000. Statistics presented in the qualitative section are unweighted.
@@
The mean respondent age was 46.0 years ($sd=16.7$). Women accounted for 50.9% of the respondents, men for 47.6%, and together “other,” “prefer not to say,” and “non-binary” accounted for the remaining 1.4%. Detailed demographic breakdowns by country are available in Table A1 in the Appendix. Our gender results and statistical tests focus primarily on the users who identified as women or men, excluding the 1.0% of other users for two reasons: we are using census weights, which are often limited to binary genders, and our total sample (unweighted) of non-binary respondents was 93, which limits our ability to perform statistical tests for significance.
@@ 3.3. Measures and Analysis
All quantitative analyses were performed using R Statistical Software v4.3.2 (R Core Team, 2023). We primarily used the survey (Lumley, 2023), srvyr (Freedman Ellis and Schneider, 2023), marginaleffects (Arel-Bundock, 2024), and epitools (Aragon et al., 2020) packages for statistical analyses. When calculating within-subject mean differences, we use two-tailed one-sample weighted t-tests with bootstrapped standard errors, using 1,000 replicates (e.g., differences in criminalization attitudes across different types of behaviors). When calculating confidence intervals for proportions, we use the Rao-Scott scaled chi-squared distribution for the loglikelihood from a binomial distribution (e.g., % of respondents who had experienced victimization). When calculating gender differences for binary categorical values, Pearson’s Chi-squared tests with Rao-Scott adjustments are used to determine significance, followed by the svyglm function with a quasibinomial family to calculate risk ratios (e.g., were men more likely to report having watched deepfake pornography as compared to women, and how much more likely?). When calculating between-gender mean differences, we use two-tailed two-sample weighted Welch t-tests using bootstrapped standard errors, and 1,000 replicates (e.g., difference in how much women think sharing deepfake pornography of celebrities should be criminalized as compared to men).
When evaluating differences amongst categorical variables in the unweighted qualitative data, we use Pearson’s Chi-squared test with Yates’ continuity correction (e.g., was there a significant difference the proportion of men versus women in indicating condemnation attitudes). When examining victim-blaming attitudes by country, we use a Fisher’s exact test to accommodate the small sample size.
@@
In relation to AI-IBSA, all respondents were asked three closed-ended questions, as well as one optional open-ended question presented towards the end of the survey. They were first asked whether they were familiar with the concept of deepfake pornography using the following wording: “Realistic-looking fake porn can now be created using artificial intelligence (AI) to swap the faces of pornographic actors with other people’s faces, so that it looks like they’re in a porn video. This is sometimes referred to as ‘deepfake pornography’ as it looks very realistic and can be hard to recognize as fake or digitally created. How familiar are you with this concept?”. This question offered four possible categorical response options, as seen in Table 1. Second, respondents were asked how much they thought a series of behaviors were worthy of criminalization (behaviors can be seen in Figure 1). Response options ranged from -2 (“Definitely should not be a crime”) to 2 (“Definitely should be a crime”), with the midpoint of 0 being “Not sure.” And third, respondents were asked how much they agreed or disagreed with the following statement: “People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ‘deepfake’) of them without their permission.” Respondents answered this question on a 7-point Likert scale, anchored at 1=“Strongly disagree” and 7=“Strongly agree” (we present the mean alongside the median in the results - see below) (Sullivan and Artino Jr, 2013).
@@
The remaining close-ended questions were asked of a subset of respondents, depending on their knowledge and experiences. A subset of respondents who indicated at least some familiarity (i.e., “I know a little bit about it” or “I am quite familiar with this”) were asked follow-up questions around AI-IBSA behaviors as seen in Figure 2. As part of a broader set of questions regarding the victimization and perpetration of image-based sexual abuse since the age of 18, we also asked the subset of respondents who indicated a previous experience with digitally altered images whether the relevant content was a video, and if so, whether the video was a deepfake (allowing respondents to select “Yes”, “No”, “Don’t know”, or “Prefer not to answer”). For victimization questions, the wording was intentionally not blame-casting. Similarly, perpetration questions were phrased neutrally to focus on whether the respondent had ever done an action.
@@
All respondents could opt into answering the open-ended question with the following wording: “Would you like to add any additional thoughts on the topic of deepfakes? Please feel free to skip this question if you don’t have anything to add.” We analyzed the qualitative data for the open-ended response questions using an iterative, inductive approach. Our approach was to translate the responses from origin language to English via the Google Translate API, familiarize ourselves with the data, develop a codebook, code the data, and then review the data to make sense of the themes from the first coding process. To do so, two members of the research team read through 25% of the sample responses across the countries and then co-developed a codebook that detailed descriptions and restrictions on what could be included within a code and provided concrete examples of each code. The same two members of the research team coded the first 150 responses based on four preliminary codes that emerged from the initial review of the data. Coder 1 and coder 2 independently moved through the subset of data labelling the open-ended responses with relevant codes. Following this, coders 1 and 2 compared coding allocations for quality assurance. The two coders then identified and removed responses that were deemed inappropriate or irrelevant to the research (e.g., “no comment,” or “no thanks, I do not know this topic”). Finally, the team revised and discussed the themes that emerged from the responses (n = 2,474). The final codebook contained four key themes: “condemnation,” “punitive and justice attitudes,” “a fear of technology,” and “victim-blaming.” Inter-rater reliability (IRR) for each code was calculated using Cohen’s Kappa. Across the four codes, Kappa values ranged from .135 to .225, with a median of .173.
@@ 4. Results
Results are organized into four sections. The first section addresses the first research question, focusing on awareness and attitude questions that were asked of all respondents. The second section answers the second question by focusing on behaviors involving AI-IBSA/deepfake pornography, including victimization, perpetration, and consumption. The third section investigates the role of gender in influencing answers to the first and second research questions. Finally, the last section is a qualitative analysis of the open-ended opinion question. We present our results primarily descriptively and holistically, particularly at the country level, as we have no reason to assume differences between countries, given the lack of prior data. When discussing gender differences, statistical analyses (accounting for weights) are presented, alongside measures of uncertainty. We conclude with some qualitative trends from the single, open-end question to which respondents could choose to respond.
@@ 4.1. Awareness and Attitudes
. Country Response % of 95% CI Respondents Australia I have never heard of this until now 36.8 35.4-38.2 I have only heard the phrase 30.3 29-31.7 I know a little bit about it 24.7 23.4-26 I am quite familiar with this 8.2 7.3-9.1 Belgium I have never heard of this until now 40.1 39.1-41.1 I have only heard the phrase 28.0 27.1-28.9 I know a little bit about it 25.6 24.7-26.5 I am quite familiar with this 6.3 5.8-6.8 Denmark I have never heard of this until now 34.8 34.1-35.5 I have only heard the phrase 44.1 43.4-44.8 I know a little bit about it 17.0 16.5-17.6 I am quite familiar with this 4.10 3.8-4.4 France I have never heard of this until now 55.2 52.8-57.6 I have only heard the phrase 21.6 19.7-23.6 I know a little bit about it 18.4 16.6-20.3 I am quite familiar with this 4.8 3.8-5.9 Mexico I have never heard of this until now 51.7 48.8-54.7 I have only heard the phrase 22.3 19.9-24.8 I know a little bit about it 21.0 18.7-23.5 I am quite familiar with this 4.9 3.8-6.3 Netherlands I have never heard of this until now 29.0 27.9-30.1 I have only heard the phrase 31.7 30.6-32.8 I know a little bit about it 31.8 30.7-32.9 I am quite familiar with this 7.5 6.9-8.2 Poland I have never heard of this until now 50.9 49.1-52.8 I have only heard the phrase 30.1 28.4-31.8 I know a little bit about it 15.2 13.9-16.5 I am quite familiar with this 3.8 3.2-4.6 South Korea I have never heard of this until now 31.1 29.1-33.2 I have only heard the phrase 39.2 37.1-41.4 I know a little bit about it 26.6 24.7-28.6 I am quite familiar with this 3.0 2.3-3.8 Spain I have never heard of this until now 49.5 47.5-51.5 I have only heard the phrase 27.6 25.8-29.4 I know a little bit about it 19.5 17.9-21.1 I am quite familiar with this 3.5 2.8-4.3 USA I have never heard of this until now 42.5 37.7-47.4 I have only heard the phrase 26.9 22.7-31.4 I know a little bit about it 23.7 19.6-28.1 I am quite familiar with this 6.9 4.5-9.8
@@
Figure 1. Heatmap of Mean Criminalization Attitudes by Country. Scale ranges from -2 (“Definitely should not be a crime”) to 2 (“Definitely should be a crime”), with a midpoint of 0 “Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parentheses.
@@
[Heat map with country on the x axis and experience type on the y axis]Heat map with country on the x axis and experience type on the y axis. Mean weighted scores are presented alongside standard errors in parentheses. All scores are positive, indicating that on average, respondents fell on the side of “should be a crime.”
@@
Figure 1. Heatmap of Mean Criminalization Attitudes by Country. Scale ranges from -2 (“Definitely should not be a crime”) to 2 (“Definitely should be a crime”), with a midpoint of 0 “Not sure”). Brighter red is more deserving of criminalization, darker blue is less deserving of criminalization. Standard errors are presented in parentheses.
Because all respondents in the survey received some information about deepfake pornography through the phrasing of the awareness question, respondents were then asked whether they thought a series of actions should be a crime, with response options ranging from -2=“Definitely should not be a crime” to 2=“Definitely should be a crime.” The midpoint of 0 was a “Not Sure” option. Weighted means (Sullivan and Artino Jr, 2013) and standard errors by country and behavior are presented in Figure 1 (with brighter red indicating more deserving of criminalization and darker blue less deserving), and weighted medians can be found in Figure A3 in the Appendix. We intentionally included “traditional” forms of image-based sexual abuse (e.g., the taking, creating, threatening to share, or sharing of real non-consensual content) to better understand how users perceived AI-IBSA/deepfake pornography in comparison. Aggregated across countries, where equivalencies existed, respondents generally rated actions involving non-consensual real content to be worse than the equivalent of AI-generated content. For example, posting/sending of real content without consent was seen as significantly more deserving of criminalization than sharing deepfake pornography of ordinary people ($M_{difference}=0.27,t=38.76,SE=0.01,p<0.001$, Cohen’s $d\ =0.30)$. Non-consensually filming or photographing someone was seen as significantly more deserving of criminalization as compared to making deepfake pornography (“ordinary people” for both, $M_{difference}=0.56,t=63.51,SE=0.01,p<0.001$, Cohen’s $d\ =0.50$).
We additionally disaggregated the deepfake pornography categories by victim-type; that is, celebrities versus non-celebrities. We hypothesized that respondents would have more lenient views towards deepfake pornography behaviors involving celebrities, who are more commonly victimized than non-famous individuals. The difference between paying someone to create deepfake pornography of a celebrity versus an ordinary person was statistically significant ($M_{difference}=0.03,t=5.01,p<0.001,SE=0.01$, Cohen’s $d\ =0.04)$, however the effect size is negligible. Similar findings can be seen for viewing deepfake pornography of ordinary people versus celebrities ($M_{difference}=0.03,t=4.63,p<0.001,SE=0.01$, Cohen’s $d\ =0.03$). This suggests little difference in perceptions of criminalization deservedness between celebrity and non-celebrity victims of equivalent behaviors.
With regard to deepfake pornography behaviors, we can draw four main conclusions. First, echoing the real content findings, the distribution of deepfake pornography was considered especially bad, as compared to something more passive such as viewing deepfake pornography (“ordinary people” for both, $M_{difference}=0.68,t=75.25,p<0.001,SE=0.01$, Cohen’s $d\ =0.60$). Second, paying someone to create deepfake pornography was considered worse than making it yourself (“ordinary people” for both, $M_{difference}=0.33,t=46.45,p<0.001,SE=0.01$, Cohen’s $d\ =0.36$). One possible interpretation is that respondents assume that going to significant lengths to have deepfake pornography created is indicative of malicious intent. Third, across countries, the behavior with the lowest average score (i.e., least bad) was the viewing of deepfake content.
@@
Finally, respondents were asked about how much they agreed or disagreed with the following statement: “People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ‘deepfake’) of them without their permission,” with anchors of 1=Strongly disagree and 7=Strongly agree. As seen in Table 4.1, across the different countries, the mean response fell between 1 and 2, with a median of 1 in all countries, indicating a general sentiment that people should get upset if they are the target of a deepfake pornography video.
@@
Table 2. Level of agreement with statement: “People shouldn’t get upset if someone creates a digitally altered video (e.g., fake porn or a ‘deepfake’) of them without their permission” (1 = strongly disagree, 4 = neither agree not disagree, 7=strongly agree). Weighted statistics presented.